You purchase a Php1,000 coupon band. The coupon rate on the bond is 4%. What is the coupon payment?

1
Expert's answer
2021-11-25T09:29:33-0500

"coupon rate= \frac {coupon payment}{bond face value} \times 100%"


"coupon payment= \frac {coupon rate}{100} \times bondfacevalue"


"coupon payment= \frac {4}{100} \times 1000"



"coupon payment= \frac {4000}{100}"


coupon repayment=40

coupon repayment=Php 40